# TerraMind: Large-Scale Generative Multimodality for Earth Observation

### v3

@@ Abstract
We present TerraMind, the first any-to-any generative, multimodal foundation model for Earth observation (EO). Unlike other multimodal models, TerraMind is pretrained on dual-scale representations combining both token-level and pixel-level data across modalities. On a token level, TerraMind encodes high-level contextual information to learn cross-modal relationships, while on a pixel level, TerraMind leverages fine-grained representations to capture critical spatial nuances. We pretrained TerraMind on nine geospatial modalities of a global, large-scale dataset. In this paper, we demonstrate that (i) TerraMind’s dual-scale early fusion approach unlocks a range of zero-shot and few-shot applications for Earth observation, (ii) TerraMind introduces “thinking in modalities” (TiM)—the capability of generating additional artificial data during finetuning and inference to improve the model output—and (iii) TerraMind achieves beyond state-of-the-art performance in community-standard benchmarks for EO like PANGAEA. The pretraining dataset, the model weights, and our code have been open-sourced under a permissive license.
@@ 1 Introduction
Earth observation (EO) increasingly benefits from multimodality because of the important integration of complementary information from different data sources. This becomes particularly relevant as data from specific satellite missions can be unavailable for a specific time or location due to low revisiting times or weather phenomena like cloud coverage. Vice versa, for the computer vision domain, EO data is an important playground for the development of new approaches as there is significant publicly available data of very high quality and complexity. The available modalities range from sensors of different satellite missions to relevant complementary information like digital elevation.
@@
In this work, we introduce TerraMind, the first large-scale, any-to-any generative, multimodal model for EO. With TerraMind, we introduce a dual-scale pretraining on pixel-level and token-level for generative multi-modality extending previous approaches like 4M [52] and demonstrate benefits over training primarily on tokens. TerraMind encodes high-level contextual information in tokens to enable correlation learning and scaling, while, additionally capturing important fine-grained representations using pixel-level inputs. During the pretraining, TerraMind predicts masked target tokens so that our pretraining objective boils down to a cross-modal patch classification problem that results in high-quality latent spaces. TerraMind is pretrained on a custom global-scale geospatial dataset named TerraMesh with 9 million samples that have been aligned spatiotemporally and across modalities [7]. In addition to radar and optical satellite imagery from the Copernicus Sentinel-1 (S-1) and Sentinel-2 (S-2) missions, our dataset contains task-specific modalities such as land use/land cover (LULC) and normalized difference vegetation index (NDVI) maps, metadata like digital elevation models (DEM) and geographic coordinates, and natural language in the form of captions. To the best of our knowledge, TerraMind represents the first truly generative, multimodal deep learning model for EO. Furthermore, in contrast to many current remote sensing models that utilize masked autoencoders (MAE) like [54], contrastive learning (CL), or diffusion techniques, TerraMind uniquely demonstrates benefits of leveraging the autoregressive paradigm for Earth observation.
@@
We provide an overview on TerraMind’s performance in the community-standard benchmark PANGAEA [49] in Figure 2 and highlight the generative capabilities of TerraMind in Figure 3. In summary, our key contributions are as follows: (i) We introduce a dual-scale approach for generative multimodal pre-training leveraging data on pixel-level and token-level, which outperforms other fusion approaches, enables generation of artificial data, and facilitates zero-shot and few-shot applications. (ii) We introduce thinking in modalities – similar to chain-of-thought approaches in LLMs – for multi-modal models in EO, demonstrating that infusing generated data during finetuning improves the performance. (iii) We demonstrate that TerraMind outperforms other geospatial foundation models both in unimodal and multimodal settings.
@@
Figure 2: TerraMind outperforms other geospatial foundation models on PANGAEA benchmark [49] in finetuning. Performance is measured in mIoU and min-max scaled per dataset with max performance displayed in parentheses.
@@
Figure 3: Chained generation example of TerraMindv1-B starting from either optical, radar, or digital elevation data. Left is input, middle is artificially generated data by TerraMind, right represents the sensor measurements and ground truths, respectively. Note that TerraMind performance is upper-bounded by the tokenizer performance, especially with respect to the diffusion decoder.
@@ 2 Related Work
Computer vision in Earth observation. Computer vision (CV) has significantly advanced Earth observation (EO) [76]. Many CV techniques, originally developed for natural image processing, have been adapted to EO [62], often with minimal modifications. A wide range of tasks benefit from these methods, including classification [16], semantic segmentation [72] (e.g., land cover mapping [20, 21]), change detection [59] (e.g., disaster response [19]), object detection [39] (e.g., vessel identification [55]), and regression (e.g., biomass estimation [53]). Deep learning architectures like CNNs [75] and Vision Transformers (ViTs) [17] have demonstrated strong performance, often surpassing traditional remote sensing (RS) methods. However, EO presents unique challenges, including diverse sensor modalities [4] and geospatial heterogeneity [46]. An emerging paradigm in EO is self-supervised learning (SSL) [64] and geospatial foundation models (GFMs) [45], which aim to leverage vast amounts of unlabeled RS data to develop general purpose task models [32]. While off-the-shelf CV models have shown promising results [36], they do not fully exploit the unique characteristics of geospatial data. Many GFMs still rely on generic CV architectures [50], which were not explicitly designed to handle the complexities of EO, such as heterogeneous sensor sources (e.g., optical, radar, DEM) [29], integrated with auxiliary data (e.g., text) [42, 47], and expert knowledge (e.g., prioritizing specific bands or indexes). In this direction, TerraMind better integrates domain-specific properties, developing a customized and expandable multimodal learning strategy.
@@
Multimodality in CV. The current paradigm of multimodal CV is driven by the integration of diverse data streams [69], such as natural images [74], natural language text [10], temporal video data [58], and weather [70], within large foundation models [8]. Starting from the alignment of images and texts [57], these models moved beyond simple feature extraction, towards nuanced contextual understanding. The ability to combine several modalities allows for unprecedented capabilities in complex tasks [30], evidenced by the rapid advancement of multimodal Large Language Models (MLLMs) [30], that excel in tasks such as scene understanding [12], visual question answering [18], and video analysis [24]. Recent advances in architectural design [31] and large scale pre-training [11] have enabled the development of models that learn highly effective cross-modal representations [41], which can then be adapted to a wide variety of downstream tasks [66].
@@
Multimodality in EO. Multimodality in EO originates from data fusion and is typically understood as the integration of SAR and optical data [13, 28, 38, 25] or the combination of optical data with vector data [5]. Some studies have explored alternative combinations of data. In [15], the authors introduce a contrastive framework for comparing RS images and street views. Even different optical sensors can be considered different modalities [61, 48]. Similarly, several multi-view images (i.e. multimodal) datasets [44, 26, 54] are introduced. More recent approaches combined text and images [40], both for discriminative [42] and generative [34] purposes. Lately, different GFMs are trained in a multimodal way [4, 54, 68], still focusing either on a specific set of modalities (e.g., vision [68], [3]) or tasks (e.g., generative [34]). Compared to multi-scale high-quality generation models for optical data, like MetaEarth [71], our approach allows to generate any modality from any other pretraining modality. To the best of our knowledge, no existing model has combined a wide and diverse amount of modalities both for discriminative and generative purposes, as TerraMind does. In Table 1, we provide a comparison of TerraMind with related model architectures.
@@
Table 1: Comparison of TerraMind to other model architectures. While several model architectures have been developed to process multiple EO modalities, TerraMind represents the first truly generative model that can produce any modality from any other modality.
@@ 3 Dataset
For the pretraining of TerraMind and its tokenizers, we create a multimodal dataset called TerraMesh, which will be open-sourced to the community. TerraMesh builds on existing datasets, which we expand by adding modalities from external data sources or by applying pseudo-labeling. We provide an overview of the aligned image modalities and a detailed dataset description in the supplementary material.
Base datasets. TerraMesh is based on SSL4EO-S12 [65, 6] and MajorTOM-Core [23], two unlabeled remote sensing datasets containing co-aligned radar and optical imagery from Sentinel-1 and Sentinel-2 satellites. SSL4EO-S12 has lower geographic coverage but is multi-seasonal. MajorTOM-Core covers most of the Earth’s land surface at a single timestamp. For MajorTOM-Core, we apply a subsampling scheme based on LULC classes and ecoregions. TerraMesh includes a total of approximately 9 million globally distributed samples from both Sentinel-1 and Sentinel-2, each measuring 264$\times$264 pixels at 10m resolution.
Additional modalities. We obtain co-aligned yearly LULC maps by ESRIhttps://planetarycomputer.microsoft.com/dataset/io-lulc-annual-v02 with nine land use classes. Additionally, we leverage SEnSeI v2 [22] as a cloud and ice annotation model and update the ESRI LULC classes for better spatiotemporal alignment. NDVI maps are computed using the corresponding spectral bands from Sentinel-2. DEM is extracted from the Copernicus DEM 30m dataset [2], which provides global coverage of the Earth’s elevation at a 30m resolution. Captions are generated synthetically by constructing RGB images from Sentinel-2 patches using the corresponding spectral bands and processing them with LLaVA-Next [37]. A tailored prompt guides the model to describe the content of each image (see supplementary material for details). For geolocations, we round the latitude and longitude values from the center of each patch to the nearest quarter degree and save the resulting discretized coordinates as strings in a well-defined format (e.g., “lat=35.75 lon=-89.00”).
@@ 4 Methods
TerraMind pretraining is two-staged. We first pretrain unimodal tokenizer models, tokenize the modalities, and then leverage token-level and pixel-level input to pretrain the TerraMind encoder-decoder architecture. In the following, we describe those individual stages.
@@ 4.1 Tokenization
We develop modality-specific tokenizers to encode each modality into a sequence of discrete tokens for pretraining and decode token sequences back to their original forms. Thus, TerraMind is in principle compatible with any modality, as long as it can be tokenized and aligned with other modalities. For reasons of space, we delegate most experiments related to the tokenizer performances to the supplementary material.
Image-like modalities. We train autoencoder-based architectures with a quantization step in the bottleneck for image-like modalities such as Sentinel-1, Sentinel-2, LULC, NDVI, and DEM. Tokenizer encoders process an input image and generate a latent representation for each 16$\times$16 patch, which is then discretized with finite-scalar-quantization (FSQ) [51] into one of $N$ codewords, where $N$ represents the codebook size. All tokenizers use a vocabulary size of 16K besides the simpler LULC modality for which we use 4K. These codewords are then used by the diffusion decoder to reconstruct the original image. The benefit of leveraging diffusion decoders lies in facilitating cross-modal generation in TerraMind by transforming tokens back into images. By mapping each codeword to a unique integer in $\{0,1,\dots,N-1\}$, we obtain discrete tokens for each image patch. We pretrain the tokenizers in a self-supervised setting. FSQ as quantization method enhances training stability [51] compared to vector quantization [63]. FSQ improves training stability by using a fixed codebook instead of a learnable one, eliminating the need for codebook-related loss terms. Notably, FSQ is heavily influenced by ideas of neural compression [27]. For our 12 channel Sentinel2-L2A data, we achieve compression rates of over 3000x by quantizing the data. We summarize the architecture of our tokenizers in Figure 4. The main objective of the overall tokenizer is to encode image patches consistently into discrete tokens based on semantic similarity to enable cross-modal correlation learning. Thus, the loss of certain details is an expected trade-off, since the focus is on grouping similar patches rather than preserving all fine-grained features. Naturally, more accurate reconstructions faciliate the cross-modal generation, which is an important component of TerraMind – the main focus however lies on consistent cross-modal correlation learning. We provided further details on the pretraining of the tokenizers in the supplementary material and open-source code and weights for the tokenizer models.
@@
Sequence-like modalities. We treat both captions and geolocations as text and use a single text tokenizer to process both modalities. By discretizing the geographic coordinates and representing them as strings in a structured format, we introduce special coordinate tokens into the tokenizer’s vocabulary. This approach allows us to encode geolocations as a sequence of discrete tokens, beginning with a latitude token followed by a longitude token. For the textual data, we modify the existing WordPiece tokenizer from BERT [33] rather than fitting a new tokenizer specifically for our captions since satellite imagery and corresponding captions have far less variations than natural images.
@@ 4.2 Pre-training
Architecture. TerraMind uses a symmetric Transformer-based encoder-decoder architecture proposed in [52], which is designed to process sequences of multimodal tokens. In addition to discrete tokens, TerraMind accepts pixel-level inputs, specifically satellite imagery and digital elevation maps. For pixel-level inputs, we apply learnable patch-wise linear projections to generate patch embeddings for each 16$\times$16 patch, similar to the approach used in ViT [17].
@@
Dual-scale early fusion. In contrast to [52], we not only embed token-level data but additionally leverage pixel-level data across a range of input modalities to introduce a dual-scale feature representation to support the structuring of the embedding space. Both tokens and patches represent a 16x16 pixel area. Tokens represent this area via a single discrete integer value, while the image patches describe the same area with the actual floating point sensor data. During pretraining, the model, thus, not only learns a correlation between modalities (i.e., cross-modal learning) but also between different levels of abstraction within the same modality. The low-level token information enables the cross-modal correlation learning, while adding pixel level input accounts for spatial nuances. Based on dual-scale features the model further learns to better structure pixel-level data in the embedding space via the corresponding information from the discrete token. We illustrate the pretraining paradigm in Figure 5. The model is agnostic to processing tokens or patches in the input space, while the target is generally token-level data. We use six pixel-level modalities and eight token-level modalities. Optical, radar, and elevation data is present both in pixel-space and token-space.
@@
Figure 5: Illustration of the pre-training task. Given a multimodal sample, random subsets of patches, input tokens, and target tokens are sampled from a Dirichlet distribution. Based on the encoded input, the decoder predicts target tokens for the masked input.
@@
Masking strategy. TerraMind applies a masked modeling approach in the token space following [52]. The model leverages a set of randomly selected target tokens that have to be reconstructed from a randomly selected set of input tokens and pixel-level data. During pre-training, we sample input and target data from a Dirichlet distribution:
$f(\mathbf{x}\mid\boldsymbol{\alpha})=\frac{1}{B(\boldsymbol{\alpha})}\prod_{i=1}^{M}p_{i}^{\alpha_{i}-1}$ (1)
where $\mathbf{p}=(p_{1},p_{2},\dots,p_{M})$ is a vector of probabilities, and $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\dots,\alpha_{M})$ is a vector of weighting parameters that we leverage during pretraining as a steering wheel to guide the model. $B(\boldsymbol{\alpha})$ refers to a multivariate Beta function.
@@
We opt for masked token reconstruction to familiarize the model with the absence of entire modalities, which is crucial for a high usability of a multimodal model in Earth observation. During pre-training, the model learns an internal representation of unseen modalities which is expected to benefit a range of downstream applications. In addition, sampling input and target tokens improves the computational efficiency of the pre-training, as each token is a compressed representation of a patch with compression factors of between 250x and 3000x depending on the modality. Finally, without tokenized representations of the image-like modalities, it is challenging to learn the correlation to sequence-like modalities. The overall training objective of TerraMind boils down to a cross-modal patch-level classification problem that we represent via a cross entropy loss:
@@
where $y_{i}$ is the one-hot encoded true class of token $i$, $p_{i}$ is the predicted probability for token $i$, $N$ is the total number of possible tokens. Interestingly, we can infer an upper bound loss for a random model. When a model outputs arbitrary target tokens, the cross entropy loss will collapse to the natural logarithm of the vocabulary size:
$\mathcal{L}_{\text{CE, random}}=-\sum_{i=1}^{N}y_{i}\log\left(\frac{1}{N}\right)=\log(N),$ (3)
which we validated empirically as an upper bound for the TerraMind pretraining loss. For a vocabulary size of 16K, a random model would have a validation loss of 9.68.

### v4

# _supplementary

@@ Abstract
We present TerraMind, the first any-to-any generative, multimodal deep learning model for Earth observation (EO). Unlike other approaches, TerraMind is pretrained on dual-scale representations combining both token-level and pixel-level data across modalities. On a token level, TerraMind encodes high-level contextual information to learn cross-modal relationships, while on a pixel level, TerraMind leverages fine-grained representations to capture critical spatial nuances. In this paper, we demonstrate that (i) TerraMind achieves beyond state-of-the-art performance in community-standard benchmarks, (ii) TerraMind can leverage “thinking in modalities” (TiM)—the capability of generating additional artificial data during finetuning and inference to improve the model output—and (iii) TerraMind’s dual-scale early fusion approach results in well-structured embedding spaces. Models and code have been open-sourced at https://huggingface.co/ibm-esa-geospatialand https://github.com/ibm/terramind.
@@ 1 Introduction
Earth observation (EO) increasingly benefits from multimodality because of the important integration of complementary information from different data sources. This becomes particularly relevant as EO is spatiotemporally sparse due to low revisiting times or weather phenomena like cloud coverage. Vice versa, for computer vision, EO data is an important playground for the development of new approaches as there is significant publicly available data of very high quality and complexity. The available modalities range from sensors of different satellite missions to relevant complementary information like digital elevation.
@@
In this work, we introduce TerraMind as the first any-to-any generative multimodal model for EO. With TerraMind, we introduce a dual-scale pretraining on pixel-level and token-level and demonstrate benefits over training primarily on tokens. TerraMind encodes high-level contextual information in tokens to enable correlation learning and scaling, while, additionally capturing important fine-grained representations using pixel-level inputs. During pretraining, TerraMind predicts masked target tokens so that our pretraining objective boils down to a cross-modal patch classification problem that results in high-quality latent spaces. TerraMind is pretrained on a custom global-scale geospatial dataset named TerraMesh with nine million samples that have been aligned spatiotemporally and across modalities [7]. In addition to radar and optical satellite images of the Copernicus Sentinel-1 (S-1) and Sentinel-2 (S-2) missions, our dataset contains task-specific modalities such as land use/land cover (LULC) and normalized difference vegetation index (NDVI) maps, metadata like digital elevation models (DEM) and geographic coordinates, and natural language in the form of captions. To the best of our knowledge, TerraMind represents the first truly generative, multimodal deep learning model for EO. Additionally, in contrast to other recent models that utilize masked autoencoders like [54], contrastive learning, or diffusion techniques, TerraMind uniquely demonstrates benefits of leveraging token-based pretraining for EO.
@@
We provide an overview of TerraMind’s performance in a community-standard benchmark [49] in Figure 2 and highlight the any-to-any generative capabilities of TerraMind in Figure 3. Our key contributions are as follows: (i) We introduce a dual-scale approach for generative multimodal pre-training leveraging data on pixel-level and token-level, which outperforms other fusion approaches and enhances embedding space structures. (ii) We introduce thinking in modalities – similar to chain-of-thought approaches in LLMs – for multi-modal models in EO, demonstrating that infusing generated data during finetuning improves the performance. (iii) We demonstrate that TerraMind outperforms other geospatial foundation models both in unimodal and multimodal settings.
@@
Figure 2: TerraMind outperforms other geospatial foundation models on PANGAEA benchmark [49] in finetuning. Performance is measured in mIoU and min-max scaled per dataset.
@@
Figure 3: Chained generation example of TerraMindv1-B starting from either optical, radar, or digital elevation data. Left is input, middle is artificially generated data by TerraMind, right represents ground truths and tokenizer reconstructions, respectively.
@@ 2 Related Work
Computer vision in Earth observation. Computer vision (CV) has significantly advanced EO [76]. Many CV techniques, originally developed for natural image processing, have been adapted to EO [62], often with minimal modifications. A wide range of tasks benefit from these methods, including classification [16], semantic segmentation [72] (e.g., land cover mapping [20, 21]), change detection [59] (e.g., disaster response [19]), object detection [39] (e.g., vessel identification [55]), and regression (e.g., biomass estimation [53]). Deep learning architectures like CNNs [75] and Vision Transformers (ViTs) [17] have demonstrated strong performance, often surpassing traditional remote sensing (RS) methods. However, EO presents unique challenges, including diverse sensor modalities [4] and geospatial heterogeneity [46]. An emerging paradigm in EO is self-supervised learning (SSL) [64] and geospatial foundation models (GFMs) [45], which aim to leverage vast amounts of unlabeled RS data to develop general purpose task models [32]. While off-the-shelf CV models have shown promising results [36], they do not fully exploit the unique characteristics of geospatial data. Many GFMs still rely on generic CV architectures [50], which were not explicitly designed to handle the complexities of EO, such as heterogeneous sensor sources (e.g., optical, radar, DEM) [29], integrated with auxiliary data (e.g., text) [42, 47], and expert knowledge (e.g., prioritizing specific bands or indexes). In this direction, TerraMind better integrates domain-specific properties, developing a customized and expandable multimodal learning strategy.
@@
Multimodality in CV. Multimodal CV is driven by the integration of diverse data streams [69], such as natural images [74], natural language text [10], temporal video data [58], and weather [70], within large foundation models [8]. Starting from the alignment of images and texts [57], these models moved beyond simple feature extraction, towards nuanced contextual understanding. The ability to combine several modalities allows for unprecedented capabilities in complex tasks [30], evidenced by the rapid advancement of multimodal Large Language Models (MLLMs) [30], that excel in tasks such as scene understanding [12], visual question answering [18], and video analysis [24]. Recent advances in architectures [31] and large scale pre-training [11] have enabled the development of models that learn highly effective cross-modal representations [41], which can then be adapted to a wide variety of downstream tasks [66].
@@
Multimodality in EO. Multimodality in EO originates from data fusion and is typically understood as the integration of SAR and optical data [13, 28, 38, 25] or the combination of optical data with vector data [5]. Some studies have explored alternative combinations of data. In [15], the authors introduce a contrastive framework for comparing RS images and street views. Even different optical sensors can be considered different modalities [61, 48]. Similarly, several multi-view images (i.e. multimodal) datasets [44, 26, 54] are introduced. More recent approaches combined text and images [40], both for discriminative [42] and generative [34] purposes. Lately, different GFMs are trained in a multimodal way [4, 54, 68], still focusing either on a specific set of modalities (e.g., vision [68], [3]) or tasks (e.g., generative [34]). Compared to multi-scale high-quality generation models for optical data, like MetaEarth [71], our approach allows to generate any modality from any other pretraining modality. To the best of our knowledge, no existing model has combined a wide and diverse amount of modalities both for discriminative and generative purposes, as TerraMind does. We provide a comparison in Table 1.
@@
Table 1: Comparison of TerraMind to other model architectures. TerraMind represents a first-of-its-kind generative, multimodal model.
@@ 3 Dataset
For the pretraining of TerraMind and its tokenizers, we create a multimodal dataset called TerraMesh [7], which will be open-sourced to the community. TerraMesh builds on existing datasets, which we expand by adding modalities from external data sources or by applying pseudo-labeling. We provide an overview of the aligned image modalities and a detailed dataset description in the supplementary material.
Base datasets. TerraMesh is based on SSL4EO-S12 [65, 6] and MajorTOM-Core [23], two unlabeled remote sensing datasets containing co-aligned radar and optical imagery from Sentinel-1 and Sentinel-2 satellites. SSL4EO-S12 has lower geographic coverage but is multi-seasonal. MajorTOM-Core covers most of the Earth’s land surface at a single timestamp. For MajorTOM-Core, we apply a subsampling scheme based on LULC classes and ecoregions. TerraMesh includes a total of approximately 9 million globally distributed samples from both Sentinel-1 and Sentinel-2, each measuring $\times264$ pixels at 10m resolution.
Additional modalities. We obtain co-aligned yearly LULC maps by ESRIhttps://planetarycomputer.microsoft.com/dataset/io-lulc-annual-v02 with nine land use classes. Additionally, we leverage SEnSeI v2 [22] as a cloud and ice annotation model and update the ESRI LULC classes for better spatiotemporal alignment. NDVI maps are computed using the corresponding spectral bands from Sentinel-2. DEM is extracted from the Copernicus DEM 30m dataset [2], which provides global coverage of the Earth’s elevation at a 30m resolution. Captions are generated synthetically by constructing RGB images from Sentinel-2 patches using the corresponding spectral bands and processing them with LLaVA-Next [37]. A tailored prompt guides the model to describe the content of each image as described in [47]. For geolocations, we round latitude and longitude from the center of each patch to the nearest quarter degree and store the discretized coordinates as strings in a pre-defined format.
@@ 4 Methods
TerraMind pretraining is two-staged following [52]. We first pretrain unimodal tokenizer models, tokenize the modalities, and then leverage token-level and pixel-level input to pretrain the TerraMind encoder-decoder architecture. We describe those individual stages in the following.
@@ 4.1 Tokenization
We develop modality-specific tokenizers to encode each modality into a sequence of discrete tokens for pretraining and decode token sequences back to images. Thus, TerraMind is in principle compatible with any modality, as long as it can be tokenized and aligned with other modalities. For reasons of space, we delegate most experiments related to the tokenizer performances to the supplementary material.
Image-like modalities. We train autoencoder-based architectures with a quantization step in the bottleneck for image-like modalities such as S-1, S-2, LULC, NDVI, and DEM. Tokenizer encoders process an input image and generate a latent representation for each $\times16$ patch, which is then discretized with finite-scalar-quantization (FSQ) [51] into one of $N$ codewords. All tokenizers use a vocabulary size of 16K besides the simpler LULC modality for which we use 4K. These codewords are then used by the diffusion decoder to reconstruct the original image. The benefit of leveraging diffusion decoders lies in facilitating cross-modal generation in TerraMind by transforming tokens back into images. By mapping each codeword to a unique integer in $\{0,1,\dots,N-1\}$, we obtain discrete tokens for each image patch. We pretrain the tokenizers in a self-supervised setting. FSQ as quantization method enhances training stability [51] compared to vector quantization [63] by eliminating the need for codebook-related loss terms. Notably, FSQ is heavily influenced by ideas of neural compression [27]. For example, on 12-band S-2 images, we achieve compression rates of over 3000x by applying quantization. We summarize the architecture of our tokenizers in Figure 4. The main objective of the overall tokenizer is to encode image patches consistently into discrete tokens based on semantic similarity to enable cross-modal correlation learning. Therefore, the loss of some details is an expected trade-off, since the focus is on grouping similar patches rather than preserving all fine-grained features. Naturally, more accurate reconstructions facilitate cross-modal generation, however the main focus of the pretraining lies on consistent cross-modal correlation learning. We provided further details on the pretraining of the tokenizers in the supplementary material.
@@
Sequence-like modalities. We treat both captions and geolocations as text and use a single text tokenizer to process both modalities. By discretizing the geographic coordinates and representing them as strings, we introduce special coordinate tokens into the vocabulary. This allows us to encode geolocations as a sequence of discrete tokens, beginning with a latitude token followed by a longitude token. For textual data, we modify the existing WordPiece tokenizer [33].
@@ 4.2 Pre-training
Architecture. TerraMind uses a symmetric Transformer-based encoder-decoder architecture proposed in [52], which is designed to process sequences of multimodal tokens. In addition to discrete tokens, TerraMind accepts pixel-level inputs, specifically satellite imagery and digital elevation maps. For pixel-level inputs, we apply learnable patch-wise linear projections to generate patch embeddings for each $\times16$ patch, similar to the approach used in ViT [17].
@@
Dual-scale early fusion. In contrast to [52], we not only embed token-level data but additionally leverage pixel-level data across a range of input modalities to introduce a dual-scale feature representation to support the structuring of the embedding space. Both tokens and patches represent a 16x16 pixel area. Tokens represent this area via a single discrete integer value, while the image patches describe the same area with the actual floating point sensor data. Thus, during pretraining, the model not only learns a correlation between modalities (i.e., cross-modal learning) but also between different levels of abstraction within the same modality. The low-level token information enables cross-modal correlation learning, while adding pixel level input accounts for spatial nuances. Based on dual-scale features the model further learns to better structure pixel-level data in the embedding space via the corresponding information from the discrete token. We illustrate the pretraining paradigm in Figure 5. The model is agnostic to processing tokens or patches in the input space, while the target is generally token-level data. We use six pixel-level modalities and eight token-level modalities.
@@
Figure 5: Illustration of the pre-training task. Given an encoded multimodal sample of random subsets of patches and input tokens, the decoder predicts target tokens for the masked input.
@@
Masking strategy. TerraMind applies a masked modeling approach in the token space following [52]. The model leverages a set of randomly selected target tokens that have to be reconstructed from a randomly selected set of input tokens and pixel-level data. During pre-training, we sample input and target data from a Dirichlet distribution.
@@
We opt for masked token reconstruction to familiarize the model with the absence of entire modalities, which is crucial for a high usability of a multimodal model in Earth observation. During pre-training, the model learns an internal representation of unseen modalities which is expected to benefit a range of downstream applications. In addition, sampling input and target tokens improves the computational efficiency of the pre-training, as each token is a compressed representation of a patch with compression factors of between 250x and 3000x depending on the modality. Finally, without tokenized representations of the image-like modalities, it is challenging to learn the correlation to sequence-like modalities. The overall training objective of TerraMind boils down to a cross-modal patch-level classification problem optimized via a cross entropy loss:
@@
where $y_{i}$ is the one-hot encoded true class of token $i$, $p_{i}$ is the predicted probability for token $i$, NN is the total number of possible tokens. Interestingly, we can infer an upper bound loss for a random model where the cross entropy loss will collapse to the natural logarithm of the vocabulary size $\mathcal{L}_{\text{CE, random}}=-\sum_{i=1}^{N}y_{i}\log\left(\frac{1}{N}\right)=\log(N)$.